Dr. David Bryan Carpenter

Northeast Parallel Architectures Center 111 College Place Syracuse University Syracuse, NY 13244-4100 Phone: (315) 443-5068 FAX : (315) 443-1973 email: dbc@npac.syr.edu

April 14, 2000

Recommendation for Saleh Elmohamed

Dear Colleague,

I met Saleh Elmohamed four years ago when I joined Northeast Parallel Architecture Center as a research scientist and project leader. In the intervening time I have had frequent contact with Saleh, and had the pleasure of working with him on many occasions. He has had many productive interactions with my research group.

The activities of my group are focussed on development of libraries, compilers, and other system software for parallel computing, especially using Java. Saleh has put together interesting and useful Java applications and applets, and helped to parallelize them using our HPJava software. Saleh was also involved in design of the virtual programming lab (VPL) used as part of the distance education work of NPAC. Saleh extended the functionality of VPL and made it into a robust and user-friendly package.

These contributions aside, Saleh’s main research interests are in algorithms and heuristics for physical optimization problems such as resource allocation and scheduling. He has also been looking at shared memory and message-passing programming models for parallel and distributed computing in general.

My experience leads me to believe that Saleh would become a highly valued colleague in any organization involved in development of computer software or hardware. I can highly recommend him for his considerable technical abilities, his constant enthusiasm, and his professional and personal good nature.

Yours sincerely,

Bryan Carpenter